OpenStudy (anonymous):

Help with derivatives. Explain only, no answer. Find the derivative of f(x) = 6x + 2 at x = 1.

OpenStudy (anonymous):

Do you know the general formula for a tangent to a function?\[ T(x) = \frac{f(x+\Delta x) - f(x)}{\Delta x} \] If yes, we can get straight to business :)

OpenStudy (anonymous):

Yes, I'm aware of the formula.

OpenStudy (anonymous):

To be correct that's not the actual tangen, it's the slope of the function at a point \(x\)

OpenStudy (anonymous):

But isn't it usually f(x)? I know it doesn't matter.

OpenStudy (jotopia34):

can't you just take the derivative and sub x=1 into that equation. Its the function value of the slope of the tangent line at x=1?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Wait, so I sub the line into the equation?

OpenStudy (anonymous):

All of the other questions are the exact same thing, exept with dif numbers.

OpenStudy (anonymous):

\[ \frac{d}{dx}f(x) = 6 \] so it doesn't really matter at which point you look.

OpenStudy (jotopia34):

if you graphed it, there would be a straight line going across at y=6

OpenStudy (jotopia34):

so, for each x, the y=1

OpenStudy (anonymous):

I'm not aware of this formula. Or is it something you just came up with?

OpenStudy (anonymous):

Which formula?

OpenStudy (jotopia34):

the d/dx just means derivative or f'(x)

OpenStudy (jotopia34):

f ' (x)

OpenStudy (anonymous):

Oh. What's the apostraphe for?

OpenStudy (anonymous):

There are some ways to write it as @jotopia34 already mentioned

OpenStudy (jotopia34):

d/dx = derivative with respect to x. the apostrophe means f Prime at x

OpenStudy (anonymous):

It means the first derivation with respect to x

OpenStudy (jotopia34):

that is the symbollish thingy for derivative of f(x). That's how I think of it

OpenStudy (anonymous):

\[ \frac{d}{dx}f(x) = \frac{df}{dx} = f'(x) = D_xf \]

OpenStudy (jotopia34):

f ' (x) = the derivative of the function f(x) in simple terms. Terms I think are easier to get than the formal definitions

OpenStudy (jotopia34):

like the captain is saying eloquently, in several, "languages" hee hee

OpenStudy (anonymous):

Okay, thanks. CAn you guys check my work?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

f ' x=8?

OpenStudy (anonymous):

IF you replace x with 1.

OpenStudy (jotopia34):

not quite in that when you take the derivative, the derivative of a constant (number with no x) is zero. So you're adding 6+0

OpenStudy (jotopia34):

is that wrong Captain?

OpenStudy (anonymous):

Oh, okay. So 2?

OpenStudy (anonymous):

General derivative of a term like \(x^n\):\[ \frac{d}{dx}(x^n)= nx^{n-1} \]

OpenStudy (jotopia34):

not quite in that when you take the derivative, the derivative of a constant (number with no x) is zero. So you're adding 6+0

OpenStudy (anonymous):

You just said that jot. :P

OpenStudy (jotopia34):

You are right with the 6 part, only you don't add 2 because the derivative of 2 = 0

OpenStudy (anonymous):

So if you take the derivative of a constant you could look at that as \(ax^0\) now try it and plug that into the fomula

OpenStudy (anonymous):

Oh, that's what you meant.

OpenStudy (jotopia34):

yes, if you're referring to my 6+0 comment

OpenStudy (anonymous):

So any constant in the original equation becomes 0?

OpenStudy (jotopia34):

yes

OpenStudy (anonymous):

\[ \frac{d}{dx}(ax^0) = 0 a x^{-1} = 0 \]

OpenStudy (anonymous):

I am btw.

OpenStudy (anonymous):

Okay. Thanks!

OpenStudy (jotopia34):

I had trouble with this too at the beginning. Formal definitions are tricky to understand.

OpenStudy (jotopia34):

I think of the constant thing as there is no x in back of it, so it is zero

OpenStudy (anonymous):

So if it was f(x)=-9/x with x=6, it would be -9/6? (I would simplify)

OpenStudy (anonymous):

-3/2

OpenStudy (anonymous):

if the x i in the denominator it's different :)

OpenStudy (anonymous):

So may rules!

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

many*

OpenStudy (jotopia34):

the x in the denominator is like saying x^-1\[x^-1\]

OpenStudy (jotopia34):

for example, instead of 2, its (-1). You do the same derivative as a positive number but don't forget to =bring the negative sign with you

OpenStudy (anonymous):

\[ \frac{d}{dx}\left( \frac{1}{x}\right)=\frac{d}{dx}\left( x^{-1}\right) = -1x^{-2} = \frac{-1}{x^2} \]

OpenStudy (jotopia34):

You have to do do prodt rule with x^-1 or the quotient rule. \[x^-1 = -1(x)^{-1-1}\]\[=-x ^{-2}\]

OpenStudy (jotopia34):

which is the same as -1/x^2\[\frac{ -1 }{ x^2 }\]

OpenStudy (jotopia34):

does any of this make sense?

OpenStudy (anonymous):

Not really.

OpenStudy (jotopia34):

okay, easy to do. part by part

OpenStudy (jotopia34):

You start with f(x)=-9/x right?

OpenStudy (jotopia34):

so next, you want to take the derivative. There are 2 ways of doing it. Do you know the product rule or the quotient rule?

OpenStudy (anonymous):

Quotient rule.

OpenStudy (jotopia34):

Great. So that rule is: (Derivative of the top x bottom), minus (the derivative of the bottom x the top) all over the (bottom part)^2 squared

OpenStudy (anonymous):

Wait, I don't get it.

OpenStudy (jotopia34):

We have in the quotient a top/bottom. Numerator function / denominator function

OpenStudy (jotopia34):

numerator =-9 denom. = x

OpenStudy (jotopia34):

numerator = top function denom = bottom function

OpenStudy (jotopia34):

so when I say derivative of the top, I mean derivative of whatever function is in the numerator. In your case it is -9

OpenStudy (jotopia34):

or whatever function is on "top"

OpenStudy (anonymous):

Oh, okay.

OpenStudy (jotopia34):

so its (deriv. top function)(bottom function with no derivative taken) - (deriv. of bottom function)(top function with no derivative taken)

OpenStudy (anonymous):

There is a general formula, in which you can plug it in:\[ \frac{d}{dx}\left(\frac{f}{g}\right) = \frac{f'g-g'f}{g^2} \] Take \(f(x) = \frac{9}{x}\) and plug it in:\[ \frac{d}{dx}\left(\frac{9}{x}\right) = \frac{0x-9}{x^2} = \frac{-9}{x^2} \]

OpenStudy (jotopia34):

the math notation looks sometimes like top=f'(x) (g(x) - g'(x) * f(x) / g(x)^2\[\frac{ 0(x) - (1)(-9) }{ x^2}\] \[=\frac{ -9 }{ x^2 }\]

OpenStudy (jotopia34):

in above, my -9 should just be 9. sorry!

OpenStudy (anonymous):

Now I'm very confused.

OpenStudy (anonymous):

i am sorry

OpenStudy (anonymous):

I think i will leave jotopia to it as it can get confusing when 2 people are talking bout the same topic. If u need anything simply message me :) And good luck with derivatives.

OpenStudy (jotopia34):

what works for me, is if you ask a step by step question, or say if something doesn't make sense in the steps.

OpenStudy (jotopia34):

for example, what's the first thing you are going to do to take this derivative

OpenStudy (anonymous):

I'm not sure.

OpenStudy (jotopia34):

okay, first is to figure out, what "formula" you need to take the derivative Do you know what that is?

OpenStudy (anonymous):

No.

OpenStudy (jotopia34):

Okay, because we have a fraction, we need to us the formula that is for a function that is any type of fraction. Do u know which that is?

OpenStudy (anonymous):

No, I am completely new with derivatives with fractions.

OpenStudy (anonymous):

Very slow comp.

OpenStudy (jotopia34):

okay. To do derivative of a fraction, there is a "formula". The "formula" IS the quotient rule

OpenStudy (jotopia34):

Here is the formula:

OpenStudy (jotopia34):

dang, can't see it. ok that reads: \[f(x) =\frac{ 9 }{ x } \] this is your question

OpenStudy (jotopia34):

\[\frac{ d }{ dx } f(x) = f'(x)\] What the above is saying is the derivative (d/dx) of your function , f(x) is the same notation as f'(x)

OpenStudy (jotopia34):

All that is saying is that now we are going to take the derivative of your function 9/x

OpenStudy (jotopia34):

for the picture, the quotient rule is \[\frac{ d }{ dx }\left[ \frac{ f(x) }{ g(x) } \right]=\frac{( f'x)(gx)-(g'x)(fx) }{ (gx)^2 }\]

OpenStudy (jotopia34):

quiz time. What is the f(x) of your question

OpenStudy (anonymous):

9/x

OpenStudy (jotopia34):

actually, when you have a fraction, you have to separate it into 2 parts. top part and a bottom part. Each part gets a function name.

OpenStudy (jotopia34):

So looking at the picture of the quotient rule equation, which is on the top, f(x) or g(x)

OpenStudy (anonymous):

?

OpenStudy (jotopia34):

Does it nmake any sense that we have to divide the fraction into 2 parts

OpenStudy (anonymous):

Yes.